Walt made an extra $9000 last year from a part-time job. He invested part of the money at 9% and the rest at 8%. He made a total of $770 in interest. How much was invested at 8%? Please show all work....
Answer by mangopeeler07(462)

x=9% money
y=8% money
coefficient=percent

x+y=9000----------------------because there was $9000 extra
9x+8y=77000-------------------because 9% x + 8% y= 77000% (or 770)

Solve for x in the first equation
x=9000-y

Plug that in
9(9000-y)+8y=77000

Distribute the 9 to the parentheses
81000-9y+8y=77000

Combine like terms
81000-y=77000

Subtract 81000 from both sides
-y=-4000

Divide by -1
y=4000

y=8% money

money invested at 8%: $4000